OpenStudy (anonymous):

@GoldEverything You're really helpful, could you help me with this one too? Dan is making a border for a triangular placemat. He wants it to have side lengths of A=8 in. and B=4 in. If the measure of angle C= 50 degrees, how many different placemats can Dan make? a. 0 b. 1

OpenStudy (anonymous):

They're basically asking if they can make a triangle witht he given triangle hence the number of placemats.

OpenStudy (anonymous):

I think the answer is 1, is that right?

OpenStudy (anonymous):

 A triangle has 6 parts, 3 angles (the uppercase letters) and 3 sides (the lowercase letters) and anyways if we have sides a=8 and b=4 and the angle in cluded (C is the angle between sides a and b) is 50 degrees .

OpenStudy (anonymous):



OpenStudy (anonymous):

K, thanks for the triangles.

OpenStudy (anonymous):

Is the answer 1?

OpenStudy (anonymous):

I would presume, is this multiple choice?

OpenStudy (anonymous):

Yes, it is

OpenStudy (anonymous):

You would have to find the third sidee (c) with the law of cosines c² = a² + c² - 2bc cosC

OpenStudy (anonymous):

Cool, thanks for the help.

OpenStudy (anonymous):

\[c^2=4^2+8^2-2(4)(8)(\cos 50)\]

OpenStudy (goldeverything):

sorry i was so late hopefully pie helped you out really well

OpenStudy (anonymous):

I got C=6.23

OpenStudy (anonymous):

@GoldEverything, it's okay! Thanks so much for the help earlier! Pie is really helping out!

OpenStudy (anonymous):

Ok, and since you got a value for c tyou now have to try this theorem, two of the sides of the traingle have to be greater than the last side so in other words

OpenStudy (anonymous):

a=4 b=8 c=6.23 a+b>c a+c>b b+c>a

OpenStudy (anonymous):

If any of those aren't true, then you dont have a triangle and so its - and if they all works then you have 1 triangle, tell me what you get for each

OpenStudy (anonymous):

Cool. Solving it right now!

OpenStudy (anonymous):

12> 6.23, So that's true. 10.23> 8, So that's true. 14.23> 4 So there is one triangle

OpenStudy (anonymous):

I get it now, thanks for all the help.

OpenStudy (anonymous):

Yep, so that's your answer, if you enjoyed my experience with me, be sure to fan and write something while you're at it =), also just let me know if you need anymore help

OpenStudy (anonymous):

Okay, thanks again. :)

OpenStudy (goldeverything):

sorry again for being late if you need any more help just ask me or my friend @doulikepiecauseidont

OpenStudy (anonymous):

It's fine! :) Will do! You both have really helped!

OpenStudy (goldeverything):

thats the point for open study were just working my the code of conduct

OpenStudy (anonymous):

Haha yea